OpenStudy (anonymous):

Give an example of a function with both a removable and a non-removable discontinuity.

OpenStudy (amistre64):

well, what makes a discontinuity removable?

OpenStudy (anonymous):

It's one that can be fixed to create a straight line. While non-removable means it cannot be fixed.

OpenStudy (amistre64):

well, mathically ... i would say that a removable discontinuity can be canceled top to bottom ... k/k = 1 after all, and 1*a = a

OpenStudy (anonymous):

My teacher explained it by using common everyday examples and that's the only way I remembered it lol.

OpenStudy (amistre64):

so, lets for a removable setup:\[\frac{(x-r)}{(x-r)}\] something that cant be removed has no common factor to cancel it out sooo\[\frac{(x-r)}{(x-r)(x-c)}\]has both a removable and a non removable discontiuity ...

OpenStudy (anonymous):

So for removable would be like \[\frac{ x-5 }{ x+5 }\] or would it be two different numbers?

OpenStudy (amistre64):

does (x-5) cancel out (x+5) ?

OpenStudy (anonymous):

Oh okay. So it has to cancel out.

OpenStudy (amistre64):

lets chk ... say x=3 \[\frac{ 3-5 }{ 3+5 }=\frac{-2}{8}\ne1\]they dont cancel

OpenStudy (amistre64):

yeah, to remove it, you need to be able to 'remove' it .... to 'cancel' it out

OpenStudy (anonymous):

Thanks for the explanations. Really helped!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks!